OpenStudy (anonymous):

Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.) I need this in pi form.

OpenStudy (e.cociuba):

A square inscribed in a circle. s = 6 r = (1/2) (diagonal of square) diagonal of square = hypotenuse of isosceles triangle with equal sides of length 6. 6^2 + 6^2 = h^2 36 + 36 = h^2 72 = h^2 h = 6√2 r = (1/2)h r = (1/2)(6√2) r = 3√2 Area of circle: A = pi r^2 A = pi(3√2)^2 A = 18pi Area of square: A = s^2 A = 6^2 A = 36 There are four segments formed by the square and circle. Area of a segment: (area of circle - area of square)/4 (18pi - 36)/4 = (9pi - 18)/2 = (9/2)pi - 9 = 5.1 in^2

OpenStudy (anonymous):

Thanks!

OpenStudy (e.cociuba):

np:)